# Synthetic data enables context-aware bioacoustic sound event detection

### v1

@@ Abstract
We propose a methodology for training foundation models that enhances their in-context learning capabilities within the domain of bioacoustic signal processing. We use synthetically generated training data, introducing a domain-randomization-based pipeline that constructs diverse acoustic scenes with temporally strong labels. We generate over 8.8 thousand hours of strongly-labeled audio and train a query-by-example, transformer-based model to perform few-shot bioacoustic sound event detection. Our second contribution is a public benchmark of 13 diverse few-shot bioacoustics tasks. Our model outperforms previously published methods by $49\%$, and we demonstrate that this is due to both model design and data scale. We make our trained model available via an API, to provide ecologists and ethologists with a training-free tool for bioacoustic sound event detection.
@@ 1 Introduction
Foundation models such as GPT-4 [1] can learn new tasks at inference time from only a handful of labeled examples—a process known as few-shot or in-context learning (ICL) [2]. This property is particularly attractive for application-driven ML domains—like bioacoustics, ecology, and conservation—where domain experts may lack extensive ML training and large labeled datasets [3, 4, 5]. However, while there is a growing push to adapt foundation models for these fields, their success is constrained by data scarcity, which limits the scope of tasks such models can be trained to perform [4].
An example of this situation occurs in few-shot bioacoustic sound event detection (FSBSED), which attempts to provide flexible modeling for the diversity of problems that arise in bioacoustics. In this task, formalized in [6], a model is provided with a support set: an audio recording, together with annotations of the onsets and offsets of the first few events of interest. The model must predict onsets and offsets of these events in the query set, which is the remainder of the recording. This type of temporally fine-scale detection is required for many applications in animal behavior and ecology [7], however the expertise and time it takes to annotate bioacoustic events has resulted in a lack of data available for training models capable of FSBSED. Prior efforts for FSBSED largely rely on a single 22-hour training dataset described in [6], leading to lightweight models tailored to small-scale data.
Figure 2: We introduce FASD13, an expansive public benchmark for few-shot bioacoustic sound event detection, which includes 13 datasets that represent a diversity of taxa, detection targets, and recording methods. Each spectrogram represents ten seconds of audio; audio from one dataset (GS) is omitted for space. Further details are in Table 1 and the Appendix.
Given the diversity of bioacoustic audio and the extreme domain gap between different few-shot detection problems [8], one expects that increasing pre-training data and model size would improve performance. A solution to the lack of a large and diverse training dataset is to construct synthetic acoustic scenes from existing unlabeled audio. This type of data is highly scalable, grants control over data attributes, and, importantly, can be generated to come along with temporally fine-scale annotations [9]. Motivated by this, we ask: Can we train a query-by-example, high-parameter model for FSBSED using only synthetically constructed scenes (Figure 1)?
To generate data, we transform unlabeled raw audio into strongly-labeled scenes using a set of custom preprocessing steps and data augmentations to inject domain randomization into the training process. These are used to train our model, which consists of a lightweight audio encoder and an encoder-only transformer, accepts the entire example at once and directly outputs sound event detection predictions for the unlabeled audio. At inference time, the model learns to make predictions from labeled demonstration audio without fine-tuning, i.e. in-context. To encapsulate the main ideas of our method, we name it DRASDIC: Domain Randomization for Animal Sound Detection In-Context.
While we focus on FSBSED, our proposed data generation method opens the door for foundation models to perform other fine-scale acoustic tasks. As a proof-of-concept, we train a zero-shot detection model (DRASD0S) that is conditioned on only a high-level set of acoustic features such as duration and frequency bounds, which could be provided a priori by a researcher familiar with the target sound.
Prior work on FSBSED has re-used validation data from [8] at test time, because there is no public FSBSED benchmark. To fill this gap, we developed FASD13: the Fewshot Animal Sound Detection-13 benchmark. Reserved exclusively for model evaluation, FASD13 complements the validation datasets already in use. FASD13 consists of 13 datasets, each of which includes audio files and annotations of event onsets and offsets (Figure 2, Table 1). They represent a diversity of taxa (birds, insects, arachnids, mammals, amphibians), detection targets (species, call type, life stage), and recording configurations (on-body, passive acoustics, laboratory). Two of these datasets, which focus on on-body recordings of carrion crows and substrate-borne drumming of jumping spiders, have not been released publicly before. The remaining eleven we curate from publicly available sources.
We evaluate DRASDIC on FASD13 and find that it represents a $49\%$ improvement on average over previously published methods. Low-parameter models widely used in the literature are not able to take advantage of the scale of data generated by our method. Our zero-shot method DRASD0S also improves over the current standard approach. These results show that synthetic acoustic scenes can be used to pre-train audio foundation models for bioacoustics tasks which are inaccessible given current data limitations.
In summary, our contributions are: 1) To enable large-scale pretraining for fine-scale bioacoustic tasks, we propose a method for generating synthetic acoustic scenes with temporally-strong labels. 2) We introduce DRASDIC, a state-of-the-art ICL detection model trained using our generated data. 3) We create FASD13, an expansive benchmark for few-shot bioacoustic sound event detection. We make our scene generator, DRASDIC weights and inference API, and benchmark publicly available. [^1] Available upon publication
@@ 2 Related Work
2.1 Few-shot sound event detection
Few-shot sound event detection was introduced in [10], and FSBSED in [6]. In bioacoustics, detection problems are highly diverse in terms of detection targets (e.g. a given species, call type, or emotional state), and therefore methods with a fixed ontology (e.g. [11]) do not apply to many situations. Moreover, labeling event onsets and offsets and fine-tuning a model for a specific problem requires expertise and computational resources that may be unavailable to practitioners. Few-shot learning attempts to address both of these challenges. [8] note central challenges in this task include sparse vocalizations, diverse target sounds, dynamic acoustic environments, and domain generalization.
Published approaches include prototypical networks [12], representation learning [13], transductive inference [14, 15]; variations and other methods are described in technical notes submitted as part of the DCASE challenge [16]. In contrast with prior approaches, we adopt a highly parameterized, transformer-based architecture which can model temporal dependencies between sparse events using stacked attention layers. Our method differs from many prior approaches in that it does not rely on fine-tuning at inference time, which can be a potential barrier for the usability of few-shot methods by non-experts.
Previous evaluation campaigns for FSBSED have centered around the DCASE challenge [8]. This challenge historically provided public training and validation datasets, and maintained a private test dataset. Subsequent efforts have either used the public validation set for both model selection and model evaluation [8], or apparently forgone model selection [13]. Our benchmark FASD13 fills this need for public evaluation datasets. In this work, we rely heavily on the validation set from [8] for model selection.
Zero-shot learning–where tasks are specified at inference without examples–has recently emerged in bioacoustics [3, 17, 4]. However, due to data scarcity, no zero-shot model has so far demonstrated temporally fine-grained detection, nor can be prompted with acoustic features. Consequently, the de-facto approach remains the signal processing method BLED [18]. Here, we demonstrate how our synthetic data pipeline addresses this limitation to enable temporally fine-grained zero-shot detection in bioacoustics.
@@ 2.1 Few-shot sound event detection
[…]
ICL refers to a model’s ability to perform a task specified through demonstrations at inference time [2]. While originally discussed in the context of language models, in the audio domain, ICL boosts speech translation [19] and text-to-speech synthesis [20]. ICL it has also been extended to fine-scale tasks in computer vision that somewhat resemble FSBSED, which include sementic segmentation [21, 22, 23] and scene understanding [24]. Similar to our method, [22, 23] employ a simple encoder-based architecture.
2.3 Synthetic data
Generative vision and audio models have been used to create data for few-shot and low-resource tasks including classification [25], detection [26], representation learning [27], and keyword spotting [28]. We are not aware of a generative audio model that produces high-quality animal sounds, and so instead developed a preprocessing pipeline to isolate potential animal sounds in publicly available data. A similar procedure was developed in [29] in a different context. For sound event detection in general audio, the cut-and-paste-based Scaper [9] library has been used to train models within a fixed ontology [30]. [31] explores the performance impacts of training on a mixture of Scaper’s synthetic acoustic scenes with real data. A challenge of training with synthetic data is that the distribution of the generated data may not match the real-world distribution [32]. Domain randomization, in which attributes of synthetic data are randomized and allowed to take on potentially non-realistic values, has been proposed as a strategy to overcome this domain gap [33]. Prior applications include robotics [34] and computer vision [35].
@@ 3.1 Data Generation
We propose a two-stage approach to generate scenes (Figure 3). Beginning with publicly available unlabeled audio, we assemble a set of background audio tracks (5.1e5 tracks, 5540 hours) and, via extensive preprocessing, a set of short clips containing events dubbed pseudo-vocalizations (or pseudovox) (5.4e6 events, 577 hours). These are pseudo-labeled through a clustering step, so that multiple similar-sounding pseudovox can be sampled at one time.
In the second stage, which we perform on-the-fly during model training, clips are randomly sampled from these collections, manipulated with data augmentations, and combined into scenes. Prior work has shown that extensive domain randomization in synthetic data generation improves transfer to real data [33], achieving the same result as strong regularization or data augmentation. Similarly, our generated scenes may be outside of what is usually recorded in the real world, due to preprocessing steps and randomness in the scene generation process. We view this as a way of increasing test-time robustness of our method.
Preprocessing
To construct our collection of pseudovox, we obtained publicly available raw audio recordings from iNaturalist [36], Animal Sound Archive [37], xeno-canto [38], Watkins [39], and WavCaps [40]. To remove background noise, we separated each recording it into four stems using BirdMixIT [41]. For each stem, we isolated potential pseudovox: segments where the amplitude envelope exceeded 25% of the maximum amplitude of the raw recording, indicating that some acoustic event was potentially occurring. There remained a high proportion of these segments that included no clear acoustic event, so we performed a quality filtering step. For this, we obtained the final activations from BirdNET [11], applied to each segment. BirdNET is a bird sound classification model whose clip-level audio representations have been shown to be useful for other bioacoustics tasks [42]. Using manual annotations of segment quality, we trained a binary linear classifier on the corresponding BirdNET embeddings. We applied this quality filter to the potential pseudovox; those that passed the filter became the final collection of pseudovox. This procedure resulted in $M=5.4e6$ pseudovox. Finally, to obtain pseudolabels for the pseudovox, we applied $k$-means clustering to the pseudovox’s BirdNET activations. We did this for $k\in K=\{\left\lfloor{M/128}\right\rfloor,\left\lfloor{M/64}\right\rfloor,\left\lfloor{M/32}\right\rfloor,\left\lfloor{M/16}\right\rfloor,\left\lfloor{M/8}\right\rfloor\},$ to obtain different levels of cluster homogeneity. For background audio, we took the raw recordings above, along with recordings from SilentCities [43], DeepShip [44], and SanctSound [45].
Scene Generation
@@
We first sample two background tracks which are overlaid on each other. We choose a clustering level $k\in K$ and two clusters $c_{T},c_{D}$ from the clusters of level $k$. We sample a random number of target pseudovox from $c_{T}$, and a random number of distractor pseudovox from $c_{D}$. We apply reverb, resampling, time flipping, and amplitude augmentations to pseudovox, and resampling augmentations to background tracks. We paste pseudovox into the background track, one-by-one, with a random time gap between pseudovox.
We maintain a binary annotation mask for the scene. This mask is initialized with zeros, and changed to ones where target pseudovox are added. Distractor pseudovox do not change the mask; they join whatever sounds are already present in the background tracks. To generate one training example, two scenes (support and query) are generated, drawing pseudovox from the same $c_{T},c_{D}$ for both. With some probability, the background tracks of the query are chosen to be different than those of the support.
@@ 3.2 Model
Using our synthetic scenes, we train our transformer-based model DRASDIC. During training the model is presented with annotated support audio and unannotated query audio, and must predict the detection labels for the query audio.
Architecture
To emphasize the contribution of the data scale provided by our scene generation method, and noting that encoder-only architectures have been used successfully for fine-scale ICL problems in computer vision [22], we adopt a straightforward but highly parametrized BERT-like architecture. This is preceded by a simple CNN spectrogram encoder. In initial experiments we used AVES [46], which was pre-trained on animal sounds, as a frozen audio encoder. The resulting models performed poorly on the DCASE validation set [8], in particular for long-duration events. Performance on the validation set increased with a CNN encoder.
In detail, for DRASDIC support and query audio is resampled to 16 kHz, concatenated, and converted to a log mel-spectrogram (256 mels, hop size 160).The CNN encoder consists of a convolutional block followed by two residual blocks [47] (kernel sizes 7, 3, 3, hidden size 64). Mean pooling in the vertical (frequency) axis is applied after each block with kernel size 2. The frequency and hidden dimensions of outputs are flattened to one dimension and mean-pooled with kernel size 2. The final frame rate is 50 Hz. The binary support label mask is max-pooled to 50 Hz, passed through a learned per-frame label encoding, and added to the CNN outputs of the support audio. This label-enriched audio representation is fed into a transformer encoder (hidden size 768, depth 12, 12 heads) with rotary position encoding [48], flash attention [49], and SwiGLU activations [50]. A final linear layer converts each output frame into detection logits.
DRASD0S modifies this approach by replacing the support audio encoder with a feature encoder (MLP with one hidden layer of size 100, output size 768). Input features are five numbers: the typical peak frequency, high frequency, low frequency, duration, and signal-to-noise ratio (SNR) of the target event. The encoded features are concatenated with the encoded query audio and passed to the transformer.
Training
Both DRASDIC and DRASD0S were randomly initialized and trained with per-frame binary cross-entropy loss on the query labels, using AdamW [51] with $(\beta_{0},\beta_{1})=(0.9,0.999)$ and weight decay $0.01$. DRASDIC is trained on support-query pairs of total duration $\operatorname{dur}_{s}+\operatorname{dur}_{q}$ seconds. Based on initial experiments, we found that a long $\operatorname{dur}_{s}$ improved validation performance. For DRASDIC we set $\operatorname{dur}_{s}=30$ and for both DRASDIC and DRASD0S we set $\operatorname{dur}_{q}=10$; see also Section 5.6.
Model, data generation, and training hyperparameters were chosen through random search. As our model selection criterion, we used average performance on the validation datasets from [8]. Inspired by [52], we applied curriculum learning to gradually increase task difficuly during training. This linearly decays the minimum pseudovox SNR from 0 dB to a minimum of -20 dB for an initial $45e$ 4 steps (DRASD0S: $43e$ 4 steps). The learning rate is linearly increased for $41e$ 4 steps to a maximum of $52e\!-\!$ 5, and then decayed to 00 after $51e$ 5 steps (cosine schedule) using batch size of 8 (DRASD0S: 4). Parameters governing data generation are provided in the Appendix. Audio features for DRASD0S are computed during training, based on the mean power spectrum of target pseudovox. Models were implemented in PyTorch [53].
@@ 4 Public Benchmark
Table 1: Details of FASD13. Additional details are in the Appendix. Datasets with a $\dagger$ are presented for the first time here. Terrestrial and underwater autonomous passive acoustic monitoring devices are abbreviated T. PAM and U. PAM, respectively.
@@
CC Carrion Crow 10 10.00 2200 On-body Spain Corvus corone + Clamator glandarius Species + Life Stage
@@
MS [60] Marmoset 10 1.67 1369 Laboratory Laboratory Callithrix jacchus Vocalization Type
@@
A collection of public FSBSED datasets was previously provided in [6, 8], but these were designated as datasets for model development (i.e. training and validation). We complement these with FASD13, a public benchmark to be used for model evaluation (Figure 2, Table 1). FASD13 consists of 13 bioacoustics datasets, each of which includes between 2 and 12 recordings. Eleven of these datasets were used from previous studies; they were chosen for their taxonomic diversity, varied recording conditions, and quality of their annotations. Two (CC and JS) are presented here for the first time. All datasets were developed alongside studies of ecology or animal behavior, and represent a range of realistic problems encountered in bioacoustics data. Details of dataset collection and preprocessing steps are in the Appendix.
We follow the data format in [6]: Each recording comes with annotations of the onsets and offsets of positive sound events, i.e. sounds coming from a predetermined category (such as a species label or call type). An $N$-shot detection system is provided with the audio up through the $N^{th}$ positive event, and must predict the onsets and offsets of positive events in the rest of the recording.
We provide two versions of each recording in FASD13: the within-file version, in which the support and query audio come from the same original recording, and the cross-file version. In the cross-file version, which is tailored to a $5$-shot problem, the support set comes from a different recording than the query set. In more detail: Each recording in the within-file version has labels associated with one class (e.g. species label). Each class appears in more than one recording. We cut each recording after $N=5$ positive events to divide the support and query sets. Within each class label, we shuffled and re-paired the support sets and query sets. Compared to the within-file version, this version represents a harder detection problem requiring greater generalization between support and query (Appendix Figure 8).

### v2

@@ Abstract
We propose a methodology for training foundation models that enhances their in-context learning capabilities within the domain of bioacoustic signal processing. We use synthetically generated training data, introducing a domain-randomization-based pipeline that constructs diverse acoustic scenes with temporally strong labels. We generate over 8.8 thousand hours of strongly-labeled audio and train a query-by-example, transformer-based model to perform few-shot bioacoustic sound event detection. Our second contribution is a public benchmark of 13 diverse few-shot bioacoustics tasks. Our model outperforms previously published methods, and improves relative to other training-free methods by $64\%$. We demonstrate that this is due to increase in model size and data scale, as well as algorithmic improvements. We make our trained model available via an API, to provide ecologists and ethologists with a training-free tool for bioacoustic sound event detection.
@@ 1 Introduction
Foundation models can learn new tasks at inference time from a few labeled examples—a process known as few-shot or in-context learning (ICL[1]. This is attractive for application-driven ML domains—like bioacoustics, ecology, and conservation—where domain experts often lack ML experience and large labeled datasets [2, 3]. Despite growing interest in adapting foundation models for these fields, data scarcity limits the tasks they can be trained to perform [3].
An example of this situation occurs in few-shot bioacoustic sound event detection (FSBSED), which attempts to provide flexible modeling for the diversity of problems that arise in bioacoustics. In this task, formalized in [4], a model receives a support set: an audio recording with onset and offset annotations for the first few events of interest. The model must predict onsets and offsets of these events in the query set, which is the remainder of the recording.
Temporally fine-scale detection is crucial for many applications in animal behavior and ecology [5], but the time and expertise needed to annotate bioacoustic events has resulted in a lack of data available for training models capable of FSBSED. Prior efforts for FSBSED largely rely on a single 22-hour training dataset described in [4], leading to lightweight models tailored to small-scale data.
In this work, we investigate simultaneously scaling model parameters and training data volume, for a FSBSED model tailored for ICL (Figure 1). To overcome limited annotated data, we turn to synthetic data, transforming raw audio into strongly labeled scenes via custom preprocessing and augmentations that introduce domain randomization. Because our focus is ICL, we use a transformer-based few-shot model that attends to support and query audio jointly—common in ICL but rare in FSBSED. We call our method DRASDIC: Domain Randomization for Animal Sound Detection In-Context.
To evaluate performance, we introduce a new 13-dataset FSBSED benchmark, FASD13 (Fewshot Animal Sound Detection-13). DRASDIC achieves a $64\%$ average improvement over prior methods that also do not use gradient updates at inference. Ablations show improvements are due to simultaneously scaling model size, training data, and improving the few-shot mechanism. We release DRASDIC weights, inference API, and FASD13 benchmark. [^1] Available at www.github.com/earthspecies/drasdic_api.
@@ 2 Related Work
FSBSED was introduced in [4]. Challenges include sparse vocalizations, diverse target sounds, dynamic environments, and domain generalization [6]. Published methods include prototypical networks [7], representation learning [8], and transductive inference [9]. Prior evaluation of FSBSED systems has centered around the DCASE challenge [6], which provided public training and validation datasets and used a private test set. Subsequent efforts have either used the public validation set for both model selection and model evaluation [6], or skipped model selection [8].
@@
ICL refers to a model’s ability to perform a task specified through demonstrations at inference time [1]. ICL has also been extended to fine-scale tasks in computer vision that somewhat resemble FSBSED, which include sementic segmentation [10, 11] and scene understanding [12]. Similar to our method, [11] employ a simple encoder-based architecture.
Generative vision and audio models have been used to create data for few-shot and low-resource tasks including detection [13] We are not aware of a generative audio model that produces realistic and low-SNR animal sounds, and so instead developed a preprocessing pipeline to isolate potential animal sounds in publicly available data. A similar procedure was developed recently in [14] for low-resource bioacoustic classification. For sound event detection in general audio, synthetic scenes assembled from multiple clips have been used to train models with fixed [15] and open ontologies [16].
@@ 3.1 Data Generation
We propose a two-stage approach to generate scenes (Figure 2). From publicly available unlabeled audio, we derive a set of background tracks (5.1e5 tracks, 5540 hours) and a set of short clips containing events dubbed pseudo-vocalizations (pseudovox) (5.4e6 events, 577 hours); these are often animal vocalizations but may include non-biological acoustic events. These are pseudo-labeled through clustering, so multiple similar-sounding pseudovox can be sampled together.
In the second stage, performed on-the-fly during training, clips are randomly sampled from these collections, manipulated with data augmentations, and combined into scenes. Generated scenes may not always resemble real audio, due to randomness in the scene generation process. As prior work has shown that domain randomization in synthetic data generation improves transfer to real data [17], we view this as a way of increasing test-time robustness of our method.
3.1.1 Preprocessing
To construct our pseudovox set, we used public recordings from iNaturalist, Animal Sound Archive, xeno-canto222www.inaturalist.org/, www.museumfuernaturkunde.berlin/en/research/animal-sound-archive, www.xeno-canto.org, respectively, as well as Watkins [18], and WavCaps [19]. To remove background noise, we separated each recording into four stems using BirdMixIT [20]. For each stem, we isolated potential pseudovox: segments where the amplitude envelope exceeded 25% of the recording’s maximum, indicating a possible acoustic event. Many segments still lacked a clear acoustic event, so we performed a quality filtering step. We manually annotated a subset of segments for vocalization presence, then trained a binary linear classifier on the final layer BirdNET [21] activations for each segment. We applied this quality filter; passing segments became the final pseudovox set. This procedure resulted in $M=5.4e6$ pseudovox. Based on performance on a held-out test set, we estimated that 98% of these pseudovox contained a clear acoustic event. To obtain pseudolabels for the pseudovox, we applied $k$-means clustering to their BirdNET activations. We did this for $k\in K=\{\left\lfloor{M/128}\right\rfloor,\left\lfloor{M/64}\right\rfloor,\left\lfloor{M/32}\right\rfloor,\left\lfloor{M/16}\right\rfloor,\left\lfloor{M/8}\right\rfloor\},$ to obtain different levels of cluster homogeneity. We inspected a random sample of 100 clusters, rating clusters as high- or low-quality based on acoustic homogeneity. Of these, 99 were deemed high quality. For background audio, we took the raw audio above, along with audio from SilentCities [22], DeepShip [23], and SanctSound [24].
3.1.2 Scene Generation
@@
We first sample two background tracks which are overlaid on each other. We choose a clustering level $k\in K$ and two clusters $c_{T},c_{D}$ from the clusters of level $k$. We sample a random number of target pseudovox from $c_{T}$, and a random number of distractor pseudovox from $c_{D}$. We apply reverb (drawn from [25]), resampling, time flipping, and amplitude augmentations to pseudovox, and resampling augmentations to background tracks. We paste pseudovox into the background track, one-by-one, with a random time gap between pseudovox. We maintain a binary annotation mask for the scene. This mask is initialized with zeros, and changed to ones where target pseudovox are added. Distractor pseudovox do not change the mask; they join whatever sounds are already present in the background tracks. To generate one training example, two scenes (support and query) are generated, drawing pseudovox from the same $c_{T},c_{D}$ for both. With some probability, the background tracks of the query are chosen to be different than those of the support.
@@ 3.2 Model
Using our synthetic scenes, we train our model DRASDIC. During training the model is given annotated support audio and unannotated query audio, and must predict detection labels for the query audio.
3.2.1 Architecture
Noting that encoder-only architectures have been used successfully for fine-scale ICL problems in computer vision [11], we adopt a simple but highly parametrized BERT-like architecture which applies attention to support and query simultaneously. This is preceded by a CNN spectrogram encoder.
Support and query audio are resampled to 16 kHz, concatenated, and converted to a log mel-spectrogram (256 mels, hop size 160). The CNN encoder is a 2-d convolutional block and two 2-d residual blocks (ker=7, 3, 3, respectively; hidden size 64), with vertical mean pooling (ker=2) after each. Frequency and hidden dimensions are flattened and mean-pooled to a final 50 Hz frame rate. The binary support label mask is max-pooled to 50 Hz, passed to a per-frame label embedding, and added to the encoded audio. This label-enriched representation enters a transformer encoder (hidden size 768, 12 heads, 12 blocks) with rotary position encoding [26]. A final linear layer maps each frame to detection logits.
3.2.2 Training
DRASDIC was randomly initialized and trained with per-frame binary cross-entropy loss on the query labels, using AdamW [27] with $(\beta_{0},\beta_{1})=(0.9,0.999)$ and weight decay $0.01$. We used support-query pairs of total duration $\operatorname{dur}_{s}+\operatorname{dur}_{q}$ seconds. Based on initial experiments, we set $\operatorname{dur}_{s}=30$ and $\operatorname{dur}_{q}=10$.
Model, data generation, and training hyperparameters were chosen through random search. As our model selection criterion, we used average performance on the validation datasets from [6]. We applied curriculum learning to gradually increase task difficulty during training. This linearly decays the minimum pseudovox signal-to-noise ratio (SNR) from 0 dB to a minimum of -20 dB for an initial $5e4$ steps. The learning rate is linearly increased for $1e4$ steps to a maximum of 2e52e\!-\!5, and then decayed to 0 after $1e5$ steps (cosine schedule) using batch size of 8. Parameters governing data generation are provided in the GitHub repository.
@@ 4 Public Benchmark
Table 1: Details of FASD13. Datasets were chosen for their taxonomic diversity, varied recording conditions, and quality of their annotations. They were manually subsetted (prior to evaluation), to reduce computational overhead. Other (minor) preprocessing steps are described on the project GitHub. Datasets with a $\dagger$ are presented for the first time here. Terrestrial and underwater autonomous passive acoustic monitoring devices are abbreviated T. PAM and U. PAM, respectively.
@@
$C$ Carrion Crow 10 10.00 2200 On-body Spain Corvus corone + Clamator glandarius Species + Life Stage
@@
MS [34, 35] Marmoset 10 1.67 1369 Laboratory Laboratory Callithrix jacchus Vocalization Type
@@
A collection of public FSBSED datasets was previously provided in [4, 6], but were designated as datasets for model training and validation. We complement these with FASD13, a public benchmark curated for model evaluation (Table 1). FASD13 consists of 13 bioacoustics datasets, each of which includes between 2 and 12 audio files. Eleven of these datasets were derived from previous studies; they were chosen for their taxonomic diversity, varied recording conditions, and quality of their annotations. Two (CC and JS) are presented here for the first time. All datasets were developed alongside studies of ecology or animal behavior, and represent a range of realistic problems encountered in bioacoustics data. Details of dataset collection and preprocessing steps are available at the GitHub repository.
We follow the data format in [4]: Each audio file comes with annotations of the onsets and offsets of positive sound events, i.e. sounds coming from a predetermined category (such as a species label or call type). An $N$-shot detection system is provided with the audio up through the $N^{th}$ positive event, and must predict the onsets and offsets of positive events in the rest of the recording.